OpenStudy (anonymous):

Could someone please help me with this question. The points A (-2, -1, z) B (2, 4, 3) and C (10, y, -1) are collinear. Find the values of y and z.

Parth (parthkohli):

I've got a neater way to do this, if you allow. If ABC are collinear, then the area of \(\triangle ABC\) is zero.

Parth (parthkohli):

Do you know what the formula of the area of a triangle is given three points?

Parth (parthkohli):

Otherwise, just try vectors, which I think @ganeshie8 is proficient in.

ganeshie8 (ganeshie8):

that area method works too, vectors method is also nice

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

ah alright thx

ganeshie8 (ganeshie8):

AB = (4, 5, 3-z) BC = (8, y-4, -4) for AB and BC to be in same direction, the components must be proportional : \[\large \dfrac{4}{8} = \dfrac{5}{y-4} = \dfrac{3-z}{-4}\] solve y,z

ganeshie8 (ganeshie8):

let me know if that not clear enough